OpenStudy (fatimah16):

Sandra exclaims that her quadratic with a discriminant of –4 has no real solutions. Sandra then puts down her pencil and refuses to do any more work. Create an equation with a negative discriminant. Then explain to Sandra, in calm and complete sentences, how to find the solutions, even though they are not real. (10 points)

OpenStudy (mww):

This is very good question. The solutions if they are not real but be non-real complex roots. Recall that if the discriminant is <0, the quadratic function is either a positive definite or a negative definite. One such case would be \[x^2 + 4x + 5 = (x+2)^2 + 1 > 0 \] Verify that discriminant it is indeed negative: \[\Delta = b^2 - 4ac = (4)^2 - 4(1)(5) = -4 < 0\] Solving for 0 we get \[(x+2)^2 = -1 \rightarrow x+2 = \pm \sqrt{-1}\] If we define the sqrt of -1 to be iota (i) Then we can write \[x = -2 \pm i\] In general for a quadratic with \[ax^2 + bx + c = 0\] We can write the solutions using the Quadratic formula as \[x = \frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\] However if the discriminant is negative then we can further write this as \[x = \frac{ -b \pm \sqrt{-(4ac - b^2)} }{2a } = \frac{ -b \pm \sqrt{i^2(4ac-b^2)} }{ 2a }\] \[= \frac{ -b \pm i \sqrt{ 4ac - b^2}}{ 2a } = \frac{ -b \pm i \sqrt{-{\Delta} }}{ 2a }\] Thus every quadratic equation with real coefficients has at least two complex roots.